OpenStudy (anonymous):

Integration(2x+1)(x^2+x)dx... Applying the general power rule

OpenStudy (zarkon):

\[u=x^2+x\]

OpenStudy (anonymous):

What do you do with the 2x+1

OpenStudy (zarkon):

\[du=(2x+1)dx\]

OpenStudy (anonymous):

so you would just drop the du?

OpenStudy (zarkon):

\[\int(2x+1)(x^2+x)dx=\int u\,du\]